Find the standard form of the line that is perpendicular to y=3/5x+9 and contains (-5,-1) I have worked this several times and my mom has too but we are not getting one of the answers provided
answers are:
a)5x-y=5
b)5x-3y=28
c)3x+5y=5
d)5x+3y=-28
PS. this is a worksheet provided by accelerated math and not directly out of the book

Answer by Alan3354(69443)

Find the standard form of the line that is perpendicular to y=3/5x+9 and contains (-5,-1) I have worked this several times and my mom has too but we are not getting one of the answers provided
answers are:
a)5x-y=5
b)5x-3y=28
c)3x+5y=5
d)5x+3y=-28
-------------------
y=3/5x+9 is in slope-intercept form, so its slope is 3/5. The slope of a line perpendicular to it will have a slope that's the negative inverse, -5/3.
y-y1 = m*(x-x1) m = -5/3 and (x1,y1) is the point (-5,-1)
y+1 = -5/3*(x+5)
3y+3 = -5(x+5)
3y+3 = -5x - 25
5x + 3y = -28
That's d.